6/35f=18/25 solve for f
I got 21/5 but can't prove it please help

Answer by edjones(8007)

6/35f=18/25
35f/6=25/18
f=(25*6)/(18*35) multiply each side by 6/35
=150/630
=5/21
.
Proof:
35 * 5/21 = 25/3
6 / (25/3) = 6 * 3/25 = 18/25
.
Ed